The csposter Class for L^{A}TEX 2ε

Martin Reddy

April 18, 1998

1 Introduction

This document describes the csposter class (v1.1) which has been written for L^{A}TEX 2ε. This class provides a generic facility for producing posters in the University of Edinburgh corporate style. The resultant posters could be used for conference presentations, departmental posters, open day information, lecture course details, etc. You can use this class by simply starting your L^{A}TEX file with:

     \documentclass{csposter}

The default behaviour of this formatting class is to produce three-column, colour documents which are typeset as 30 point text on an A0 page. The posters can be printed on the EUCS colour A0 plotter, as well the department’s A4 laser printers.

This document will describe the various options that csposter currently supports, how to generate title sections, how to incorporate figures into posters, how to preview and print the posters, and finally, how to modify various aesthetic settings such as colour and spacing.

2 Class Options

The csposter supports the following class options which can be specified in the squarebracketed section of the \documentclass command.

• cs / lfcs : these options specify whether the poster is to be affiliated to the Department of Computer Science or the Laboratory for Foundations of Computer Science, respectively. This information is used to alter the department name and WWW address at the bottom of the poster. The default is cs.

• twocolumn / threecolumn / fourcolumn : these options enable you to change the number of columns of text used for the poster. The default is threecolumn.

• centre : this option can be used to centre the main poster text vertically on the page. The default behaviour is to present the main text directly after the title header.

• landscape / portrait : The default is landscape, using portrait re-sizes the page to be A0 portrait

3 Title Header

The title header of the poster is generated by the \maketitle command. Before issuing this command you can define a main title, a sub-title, a list of authors, and an abstract. Only the main title is required. In addition, you can provide information for an enquiries contact point.

• \title{<main-title>} : this command specifies the main title of the poster which will appear at the top of the page. This information is required.

• \subtitle{<sub-title>} : specifies the sub-title of the poster. This will appear directly below the main title.

• \author{<author-list>} : declares the authors of the poster.

• \abstract{<abstract>} : the abstract of the poster which is typeset immediately before the main text.

• \enquiries{<details>} : this can be used to provide contact information for enquiries, e.g. an e-mail or postal address. If provided, this information will replace the WWW URL at the bottom of the poster.

For example,

   \title{What I Learned in Room 101}
   \author{Winston Smith}
   \enquiries{w.smith@ministry-of-plenty.gov}
   \maketitle

4 Figures

As a result of implementing the multi-column behaviour of csposter, normal floats such as the figure and table environments are not supported. To simplify the inclusion of EPS

figures in a poster the \epsfigure[<width-percent>]{<filename>}{<caption>} command

is provided. The optional argument, <width-percent>, can be used to change the width of the EPS figure relative to the width of a single column. This is given as a number representing a percentage in the range 5..100. The default width percentage is 75%. For example:

   \epsfigure{figure.eps}{This is my figure}
   \epsfigure[100]{figure.eps}{This is a full-size version of my figure}

5 Previewing and Printing

In order to print your poster you should generate an Encapsulated PostScript (EPS) file. In order to ease this process, a Makefile has been written to perform all of the necessary operations in order to get an EPS file that you can just dump to an appropriate printer. This Makefile can be copied to the current direcory with the following shell command. See the top of the Makefile for documentation on its usage and customisation features.

   cp /usr/local/lib/TeX/tex/latex/examples/Makefile.csposter Makefile

If you then type make view, then an EPS file will be generated (e.g. poster.tex) and previewed to the screen. You can then print this file to the departmental printers by simply doing a:

     lpr -Zlandscape eps-filename

In order to print to the EUCS A0 Versatec solid-area colour plotter, then you will have to do the following:

     make scalea0
     lpr eps-filename -Peucsvc23 -Zforms=a0u

N.B. the EUCS printers can only be accessed from a certain number of servers, e.g. calvay and papa; and there is currently a limited quota for CS1-4 and M.Sc. students. Also, printing to the A0 printer is expensive, so users are encouraged to preview their posters on the departmental printers before using the A0 colour plotter.

6 Colours

Users can change the colour of text within their poster by using any of the following predefined

macros: \red, \green, \blue, \cyan, \magenta, \yellow, \black, \white, \darkgray,
\gray, \lightgray, \lightgray, or \white. For example:
   This text is black (default) and {\red this is in red}.

In addition to this, csposter displays various logical items in different colours, e.g. section headings are rendered in blue by default. The user can change the colour for any of these logical items by assigning a different colour to one of the following:

• \TitleColour : the colour used for the main title

• \SubtitleColour : the colour used for the sub-title

• \AuthorColour : the colour used for the author list

• \AbstractColour : the colour used for the abstract

• \SectionColour : the colour used for section headings

• \ItemizeColour : the colour used for itemize list bullets

• \CaptionColour : the colour used for figure captions

For example, if you want to change the default colour of the author list to be green, then you should do this as follows:

\let\AuthorColour\green

7 Spacing

A number of lengths are provided in order to customise the spacing between various components of the poster. These include:

• \SpaceAfterTitle : the extra vertical space after the main title.

• \SpaceAfterSubtitle : the extra vertical space after the sub title (if present).

• \SpaceAfterAuthor : the extra vertical space after the author list (if present).

• \SpaceAfterAbstract : the extra vertical space after the abstract (if present).

• \SpaceBeforeText : the extra vertical space between the title header and the main text of the poster.

• \SpaceToColumnSep : the horizontal space between the edge of a column’s text and the vertical line separating the columns. By default, this is set to 3em.

For example, if you want to add an extra inch of space between the title section of the poster and the main text, then you should do this as follows. Note that any change of these lengths must of course appear before the call to \maketitle.

\setlength{\SpaceBeforeText}{1in}

There is also provision for altering the balancing of columns. By default csposter will attempt to balance all columns of text so that their base lines are at the same level. You can however manually offset this balancing with the unbalance counter. This will add the specified number of lines of text to all but the right-most column. For example, \setcounter{unbalance}{5} will add an extra 5 lines of text to each column (except the last one), thus making the text appear to take up more vertical space and leave a more ragged right-hand column.

A Example L^{A}TEX 2ε Code

Above L^{A}TEX source available on-line in /usr/local/share/TeX/tex/latex/examples